### v1

@@ 3.2 Atomic Routes in Relational Deep Learning
To address the challenges outlined above, we introduce the concept of atomic routes.
@@
An atomic route is a sequence of node-types that form a composite path between the starting and ending node-type. We distinguish two cases:
@@
    Single foreign key. If a table has exactly one foreign key to another table, the atomic route is an edge connecting the primary-key node to the foreign-key node.
@@
    Multiple foreign keys. If a table has multiple foreign keys, the atomic route is a hyperedge connecting pairs of foreign-key nodes via the primary-key node.
Fig. 1 and 3 illustrates the primary-foreign key relationships in the rel-f1 dataset and the atomic routes derived from these relationships. For instance, the circuits table has only one foreign key, which points to the races table, forming atomic routes (circuits $\to$ races) and (races $\to$ circuits). In contrast, the standings table has two foreign keys connecting it to both the drivers and races tables. This results in atomic routes (drivers $\to$ standings $\to$ races) and (races $\to$ standings $\to$ drivers). These routes capture the necessary interactions among multiple entities within a single step.
Atomic routes vs. Meta-paths: Following Definition 3.1, atomic routes are constructed solely based on primary–foreign key relationships among tables, allowing them to be derived systematically without requiring expert knowledge or manual intervention. This ensures both broad applicability and scalability across diverse datasets. In contrast, meta-paths in general heterogeneous graphs often require manual specification, risk introducing bias through selective design, and may overlook crucial relational contexts. While meta-paths aim to highlight specific patterns, atomic routes serve a fundamentally different role: to systematically and comprehensively capture all essential interactions inherent in relational data.
3.3 Composite Message Passing for Relational Deep Learning
In this subsection we build upon the concept of atomic routes and design composite message passing mechanismsfor RDL. We begin this discussion by applying a standard heterogeneous GNN on a subgraph encoding tables with multiple foreign keys. We assign src, dst, and mid to represent nodes corresponding to source, destination, and intermediate node-types, respectively. In standard heterogeneous GNNs, it takes two steps to complete the full information exchange. In the first step, each mid node aggregates information from all its neighbor nodes:
@@ 3.3 Composite Message Passing for Relational Deep Learning
Note that in Eq. (1), T represents all node-types connected to the intermediate node-type. Therefore, in addition to the information from the source node-type, information from other irrelevant node-types connected to the intermediate node-type is also aggregated and entangled during this step. Specifically, information from dst is also passed to mid in this step, and it is subsequently passed back to dst again in Eq. (2), leading to redundancy, as discussed in Section 3.1.
@@
Eq. (3) describes an information exchange from src via mid to dst that is completed within a single step. As a result, there is no extraneous information entangled in the process. In summary, our approach effectively tackles the challenges that standard heterogeneous GNNs may encounter, such as multiple steps needed for complete information exchange and entangled information during message passing.
Figure 3: All the atomic routes derived from it of rel-f1 dataset. The primary-foreign key relations of rel-f1 is illustrated in Figure 1 (a).
@@ 3.4 RelGNN: Composite Message Passing with Atomic Routes
The introduction of atomic routes and composite message passing enables the design of new architectures specifically tailored to relational data graphs. Eq. (3) admits multiple instantiations, offering a flexible framework for message passing using atomic routes. Here, we propose RelGNN, a simple yet effective instantiation. When multiple foreign keys are involved, RelGNN instantiates Eq. (3) as follows: for each mid node, it fuses information from each src node connected via a primary–foreign key relationship. Then $\operatorname{FUSE}(\cdot)$ is implemented as a linear combination:
@@
Note that, due to the nature of foreign keys, each mid node is connected to only one src node. Then, RelGNN instantiates $\operatorname{AGGR(\cdot)}$ with the standard multi-head attention mechanism  (Vaswani et al., 2017; Shi et al., 2021), where embeddings from destination nodes serve as queries, and embeddings derived from the fusion operation in Eq. (4) serve as keys and values. Let $\mathbf{h}_{\texttt{fuse}}^{(l)}:=\operatorname{FUSE}(\mathbf{h}_{\texttt{mid}}^{(l)},\mathbf{h}_{\texttt{src}}^{(l)})$ as defined in Eq. 4. $\operatorname{AGGR(\cdot)}$ is realized as:
@@
In cases where tables with multiple foreign keys are not present, there is only a source and destination node-type, so the fusion operation is not needed. We directly substitute $\mathbf{h}_{\texttt{fuse}}$ with $\mathbf{h}_{\texttt{src}}$ in Eq. (3.4):
@@
We use different weight matrices in Eq. (4) and Eq. (3.4) for each atomic route, enabling RelGNN to capture different types of information across routes.
@@
where $\mathcal{T}(\texttt{dst})$ denotes the set of atomic routes with dst as the destination node, and $\mathbf{m}_{t}^{(l+1)}$ denotes the message from atomic route $t$, as defined in Eq. (3) or Eq. (6). Note that this final update operation will not lead to information entanglement, as the model learns distinct weight matrices for each atomic route, therefore learning to assign an appropriate weights to message from each atomic route during the summation.
@@ 4 Experiments
We evaluated RelGNN on RelBench (Robinson et al., 2024), a public benchmark designed for predictive tasks over relational databases using GNNs. RelBench offers a diverse collection of real-world relational databases and realistic predictive tasks. The benchmark spans 7 datasets, each carefully processed from real-world sources across a wide range of domains, including e-commerce, social networks, medical records, Q&A platform and sports. These datasets vary significantly in size, with differences in the number of rows, columns, and tables, serving as a challenging and comprehensive benchmark for RDL model evaluation. See Appendix  A.1 for the description and detailed statistics of each dataset.
RelBench introduces 30 predictive tasks covering a wide range of real-world use cases, grouped into three representative types: entity classification (Section 4.1), entity regression (Section 4.2), and recommendation (Section 4.3). These tasks are designed to reflect practical applications, such as predicting event attendance, estimating sales of an item, and recommending posts to users. The data is split temporally, with models trained on data from earlier time periods and tested on data from future time periods. The tasks vary significantly in the number of entities in the train/validation/test split and the proportion of test entities encountered during training. Description and detailed description of each task can be found in Appendix  A.2.
We follow the implementation of RDL in Robinson et al. (2024). As introduced in Sec. 2.2, relational data is transformed into heterogeneous temporal graphs, and temporal neighbor sampling is employed, creating subgraphs on which the model is trained. The initial node embeddings are extracted from raw table feature using PyTorch Frame  (Hu et al., 2024). These embeddings are subsequently fed into a GNN, and the resulting node embeddings are passed into a prediction head specific to the type of task to produce the final output. For baselines, we compare with the heterogeneous GraphSAGE  (Hamilton et al., 2017; Fey & Lenssen, 2019; Robinson et al., 2024) used in the original RelBench paper. To ensure a fair comparison, we maintain identical settings, including temporal neighbor sampling algorithm, initial node embeddings extraction model, the prediction head, and the loss function. Additionally, we incorporate a Light Gradient Boosting Machine (LightGBM)  (Ke et al., 2017) as an additional non-RDL baseline, which is applied directly to the raw entity table features, following the setting in Robinson et al. (2024).
@@ 4.1 Entity Classification
The entity classification task involves predicting binary labels for a given entity at a specific seed time. The performance is evaluated with the ROC-AUC  (Hanley & McNeil, 1983) metric, where higher values indicate better performance. The prediction head for this task consists of an MLP applied to the entity embeddings generated by the GNN. The model is trained using binary cross-entropy loss, and results are averaged over five different seeds.
Table 4.1 presents the results, along with the relative improvement of RelGNN over the standard heterogeneous GNN. RelGNN outperforms the baselines on 10 out of 12 tasks and achieves comparable performance on the remaining two. Notably, the relative gain is more significant on datasets with a more complex primary-foreign key structure (e.g., rel-f1; see Appendix B for visualizations) compared to datasets with simpler structures (e.g., rel-amazon and rel-hm). This suggests that the improvements stem from RelGNN’s ability to better capture distinct structural characteristics of relational data graphs. On the rel-stack dataset, RelGNN does not achieve significant improvements. A potential cause might be the unique self-loop structure, where primary-foreign key links connect nodes of the same type (posts)—a pattern not present in other datasets. Treating these self-loops the same way as primary-foreign key links between different node-types may not be optimal. Developing better techniques to handle self-loop structures in relational graphs is an interesting direction for future work.
@@ 4.2 Entity Regression
Entity regression task requires predicting numerical labels for an entity at a specific seed time. The evaluation metric is Mean Absolute Error (MAE), where lower values indicate better performance. Similar to the classification task, the prediction head is an MLP applied to the entity embeddings generated by the GNN, and the model is trained using L1 loss. Results are averaged over five seeds.
Table 4.2 presents the results and the relative gain of our model over the standard heterogeneous GNN. RelGNN outperforms the baselines on 8 out of 9 tasks and achieves comparable performance on the remaining one. As in the classification task, improvements are more pronounced on datasets with more complex primary-foreign key structures, highlighting RelGNN’s effectiveness in modeling relational dependencies. The performance gain on rel-stack is limited, as discussed in Section 4.1. We also observe varying performance across tasks in the rel-trial dataset. A potential cause might be the inherent limitation of the prediction head. The original RelBench paper identified issues with the entity regression prediction head, suggesting that the suboptimal performance on study-adverse may result from the prediction head rather than the GNN model itself  (Robinson et al., 2024). Notably, study-adverse requires estimating the number of severely affected patients (unbounded prediction), whereas site-success involves predicting the success rate (bounded prediction). One possible explanation is that unbounded predictions may be more challenging given the inherently constrained design of the regression head. Future work could explore alternative prediction heads better suited for entity regression tasks.
@@ 4.3 Recommendation
Recommendation tasks involve predicting a ranked list of top $K$ target entities for a given source entity at a specific seed time, where $K$ is pre-defined for each task in RelBench. This requires computing pairwise scores between source and target nodes. Following the original RelBench implementation, we employ two representative prediction heads: two-tower GNN  (Wang et al., 2019a) and identity-aware GNN (ID-GNN)  (You et al., 2021). The two-tower GNN calculates pairwise scores through the inner product of source and target node embeddings and is trained using the Bayesian Personalized Ranking loss  (Rendle et al., 2012). ID-GNN computes scores by applying an MLP to the embedding of the target entity within a subgraph sampled for each source entity and is trained using binary cross-entropy loss. Consistent with the original implementation, we use two-tower GNN for the rel-amazon dataset and ID-GNN for the remaining datasets. The evaluation metric is Mean Average Precision (MAP) @$K$, where higher values indicate better performance. Results are averaged over five seeds.
Results are presented in Table 4.3. RelGNN  achieves better or same performance compared to baselines on all 9 tasks. One limitation of the ID-GNN prediction head is that the MLP is applied to target entities sampled within the subgraph around the source node. This restricts the final recommendation list for a source node to targets that were sampled within its local subgraph. In the RDL setting, where past interactions are used to predict future ones, this restriction limits recommendations to target nodes that have previously interacted with the source node. This can be problematic in certain cases. For example, in the rel-hm dataset, the task is to predict a list of articles a customer will purchase in the next week. If candidate articles are limited to those appearing in the customer’s subgraph, the model can only recommend articles the customer has already purchased, ignoring new ones. Since customers rarely repurchase the same article, this creates an inherent limitation. This issue was identified by (Yuan et al., 2024), who proposed a quantitative metric, locality score, to measure its impact on model performance. The locality score is defined as the fraction of ground-truth target nodes that fall within the source node’s subgraph for each task. We observe that our improvements are more significant on tasks with a higher locality score (e.g., tasks in rel-stack, where users are more likely to interact with posts they have engaged with previously) compared to tasks with a lower locality score (e.g., tasks in rel-hm and rel-trial). Developing a more effective prediction head and framework for recommendation tasks is an interesting direction for future work.
@@ 5 Related Work
Several works have explored the use of GNNs for learning on relational data  (Schlichtkrull et al., 2018; Cvitkovic, 2019; Šír, 2021; Zahradník et al., 2023). These works investigated different GNN architectures that utilize the relational structure. More recently, Fey et al. (2024) introduced Relational Deep Learning (RDL) (see Sec. 2.2), establishing a new subfield of machine learning. RDL has enabled various research opportunities, such as advancements in relational graph construction algorithms, GNN architectures, and task-specific prediction heads. Yuan et al. (2024) focused on improving task-specific prediction heads for recommendation tasks, addressing limitations in the currently employed two-tower and pair-wise prediction heads. In contrast, our work focuses on improving GNN architectures applied to all task types to generate node embeddings, offering an orthogonal contribution. In addition to GNNs, Wydmuch et al. (2024) proposed leveraging large language models (LLMs) to address predictive tasks in RDL.
@@
The literature of knowledge graphs  (Bordes et al., 2013; Wang et al., 2014; 2017) differs from RDL in terms of the tasks being tackled. Knowledge graph models mainly focus on completion tasks like predicting missing entities (e.g., Q: Who is the author of ”Harry Potter”? A: J.K. Rowling) or missing relationships (Q: Did Yoshua Bengio win a Turing Award? A: Yes). In contrast, RDL focuses on making predictions about entities or groups of entities (e.g., Will a customer churn in the next month? How much will a customer spend in the upcoming week?)

### v2

@@ 4.1 Atomic Routes in RDL
To address the limitations of standard message passing in relational entity graphs outlined above, we introduce the concept of atomic routes.
@@
An atomic route is a simple path between node-types that enables a single-hop interaction between the source and destination node-type. We distinguish two cases:
@@
    Single foreign key. If a table has exactly one foreign key referencing to another table, the atomic route is an edge between the foreign key node-type (source) and the primary key node-type (destination), forming a simple path (source $\rightarrow$ destination).
@@
    Multiple foreign keys. If a table has multiple foreign keys referencing to different tables, the atomic routes are hyperedges that connect pairs of foreign key node-types (source and destination) via the intermediate primary key node-type, forming simple paths of type (source $\rightarrow$ intermediate $\rightarrow$ destination).
Figure 1 and 3 illustrate the primary-foreign key relationships in the rel-f1 dataset and the atomic routes derived from these relationships. For instance, the circuits table has only one foreign key, which points to the races table, forming atomic routes (circuits $\to$ races) and (races $\to$ circuits). In contrast, the standings table has two foreign keys connecting it to both the drivers and races tables. This results in atomic routes (drivers $\to$ standings $\to$ races) and (races $\to$ standings $\to$ drivers). These routes capture the necessary interactions among multiple entities within a single step.
Comparison to Meta-paths.
We compare atomic routes to meta-paths, a widely utilized concept in heterogeneous graph learning (cf. Section 6 for an extended discussion of meta-paths). Although both structures facilitate message propagation, they differ fundamentally in their construction and purpose. Meta-paths typically necessitate manual design guided by domain knowledge and task-specific intuition. This manual design process makes them labor-intensive, sensitive to dataset-specific characteristics, and susceptible to introducing selection bias. In contrast, atomic routes are derived automatically from the primary–foreign key relationships defined in the database schema. This enables their extraction without human supervision, ensuring applicability across a wide range of relational databases—including those with complex or previously unseen schemas. While meta-paths are often constructed to emphasize particular semantic patterns relevant to a task, atomic routes serve a fundamentally different function: they systematically capture the minimal pathways to avoid structural inefficiencies due to the primary-foreign key constraints.
4.2 Composite Message Passing for RDL
In this subsection, we build upon the concept of atomic routes and design composite message passing mechanisms for RDL. We begin this discussion by applying a standard heterogeneous GNN on a subgraph that encodes tables with multiple foreign keys. We assign src, dst, and mid to represent nodes corresponding to source, destination, and intermediate node-types, respectively. In standard heterogeneous GNNs, it takes two steps to complete the full information exchange. In the first step, each mid node aggregates information from all its neighbor nodes:
@@ 4.2 Composite Message Passing for RDL
Note that in Equation (1), T represents all node-types connected to the intermediate node-type. Therefore, in addition to the information from the source node-type, information from other node-types connected to the intermediate node-type is also aggregated during this step. Furthermore, the information from dst is passed to mid in this step, and subsequently passed back to dst again in Equation (2), leading to redundancy, as discussed in Section 3.2.
@@
Equation (3) describes a composite information exchange from src via mid to dst that is completed within a single step. As a result, there is no extraneous information entangled in the process. In summary, our approach effectively tackles the challenges that standard heterogeneous GNNs may encounter, such as multiple steps needed for complete information exchange and redundant aggregation during message passing.
Figure 3: All the atomic routes derived from rel-f1 dataset. The primary–foreign key relations of rel-f1 is illustrated in Figure 1 (a).
@@ 4.3 RelGNN: Composite Message Passing with Atomic Routes
The introduction of atomic routes and composite message passing enables the design of new architectures specifically tailored to relational entity graphs. Equation (3) admits multiple instantiations, offering a flexible framework for message passing using atomic routes. Here, we propose RelGNN, a graph attention instantiation of Equation (3). When multiple foreign keys are involved, RelGNN instantiates Equation (3) as follows: for each mid node, it fuses information from each src node connected via a primary–foreign key relationship. Specifically, $\operatorname{FUSE}(\cdot)$ is implemented as a linear combination:
@@
Note that, due to the nature of foreign keys, each mid node is connected to only one src node. Then, RelGNN instantiates $\operatorname{AGGR(\cdot)}$ with the standard multi-head attention mechanism  (Vaswani et al., 2017; Shi et al., 2021), where embeddings from destination nodes serve as queries, and embeddings derived from the fusion operation in Equation (4) serve as keys and values. Let $\mathbf{h}_{\texttt{fuse}}^{(l)}:=\operatorname{FUSE}(\mathbf{h}_{\texttt{mid}}^{(l)},\mathbf{h}_{\texttt{src}}^{(l)})$ as defined in Equation (4). $\operatorname{AGGR(\cdot)}$ is realized as:
@@
In cases where tables with multiple foreign keys are not present, there is only a source and destination node-type, so the fusion operation is not needed. We directly substitute $\mathbf{h}_{\texttt{fuse}}$ with $\mathbf{h}_{\texttt{src}}$ in Equation (4.3):
@@
We use different weight matrices in Equation (4) and Equation (4.3) for each atomic route, enabling RelGNN to capture different types of information across routes.
@@
where $\mathcal{T}(\texttt{dst})$ denotes the set of atomic routes with dst as the destination node, and $\mathbf{m}_{t}^{(l+1)}$ denotes the message from atomic route $t$, as defined in Equation (3) or Equation (6). Note that this final update operation will not lead to information entanglement, as the model learns distinct weight matrices for each atomic route, therefore learning to assign an appropriate weights to message from each atomic route during the summation.
@@ 5 Experiments
We evaluate RelGNN on RelBench (Robinson et al., 2024), a public benchmark designed for predictive tasks over relational databases using GNNs. RelBench offers a diverse collection of real-world relational databases and realistic predictive tasks. The benchmark covers 7 datasets, each carefully processed from real-world sources across diverse domains such as e-commerce, social networks, medical records, Q&A platforms, and sports. These datasets vary significantly in size, with differences in the number of rows, columns, and tables, serving as a challenging and comprehensive benchmark for RDL model evaluation. Appendix A.1 provides description and detailed statistics for each dataset.
RelBench introduces 30 predictive tasks covering a wide range of real-world use cases, grouped into three representative types: entity classification (Section 5.1), entity regression (Section 5.2), and recommendation (Section 5.3). These tasks are designed to reflect practical applications, such as predicting event attendance, estimating sales of an item, and recommending posts to users. The data is split temporally, with models trained on data from earlier time periods and tested on data from future time periods. To attach target labels, each task defines a training table that links entities of interest to their target labels and timestamps via foreign keys, enabling automatic supervision from historical data while ensuring temporal consistency during training. The tasks vary significantly in the number of entities in the train/validation/test split and the proportion of test entities encountered during training. Detailed description of each task can be found in Appendix A.2.
We follow the data processing pipeline introduced in Robinson et al. (2024). Relational data is transformed into heterogeneous temporal graphs, and temporal neighbor sampling is used to construct subgraphs centered around each entity and timestamp. Initial node embeddings are extracted from raw table attributes using PyTorch Frame (Hu et al., 2024). Final node embeddings are passed to prediction heads specific to the type of task to generate final predictions.
For baselines, we compare with the heterogeneous GraphSAGE  (Hamilton et al., 2017; Fey & Lenssen, 2019; Robinson et al., 2024) used in the original RelBench paper. To ensure a fair comparison, we maintain identical settings, including the data processing pipeline and training table, the temporal neighbor sampling algorithm, the initial node embeddings extraction model, the prediction head, and the loss function. We also incorporate a Light Gradient Boosting Machine (LightGBM)  (Ke et al., 2017) as an additional non-RDL baseline, which is applied directly to the raw entity table features, following the setting in Robinson et al. (2024).
@@ 5.1 Entity Classification
Experimental Setup.
The entity classification task involves predicting binary labels for a given entity at a specific seed time. The performance is evaluated with the ROC-AUC  (Hanley & McNeil, 1983) metric, where higher values indicate better performance. The prediction head for this task consists of a multi-layer perceptron (MLP) applied to the GNN-generated node embeddings. The model is trained using binary cross-entropy loss. All results are averaged over five different seeds.
Results.
Table 5.1 presents the results, along with the relative improvement of RelGNN over the standard heterogeneous GNN. RelGNN outperforms the baselines on 10 out of 12 tasks and achieves comparable performance on the remaining two. Notably, the relative gain is more significant on datasets with a more complex primary-foreign key structure (e.g., rel-f1; see Appendix B for visualizations). In contrast, performance improvements are smaller on datasets with simpler primary-foreign key structures, like rel-amazon and rel-hm. This trend aligns with the core design principle of RelGNN, demonstrating that its performance gains arise from effectively highlighting key predictive signals and eliminating redundant message aggregation—key challenges faced by standard heterogeneous GNNs when applied directly to relational entity graphs.
Discussion and Limitations.
On the rel-stack dataset, RelGNN does not achieve significant improvements. A potential cause might be the unique self-loop structure, where primary-foreign key links connect nodes of the same type (posts). This pattern does not present in other datasets in RelBench and introduces an orthogonal modeling challenge. We provide a potential solution in Appendix D. Since robust handling of self-loops is orthogonal to our core contribution, we leave it for future work.
@@ 5.2 Entity Regression
Experimental Setup.
Entity regression task requires predicting numerical labels for an entity at a specific seed time. The evaluation metric is Mean Absolute Error (MAE), where lower values indicate better performance. The prediction head is an MLP applied to GNN-generated node embeddings, identical to the setup used for entity classification. The model is trained using L1 loss. Results are averaged over five random seeds.
Results.
Table 5.2 presents the results and the relative gain of our model over the standard heterogeneous GNN. RelGNN outperforms the baselines on 8 out of 9 tasks and achieves comparable performance on the remaining one. As in the classification task, improvements are most pronounced on datasets with complex primary-foreign key structures, highlighting that RelGNN’s effectiveness in modeling relational structure translates to consistent gains in both classification and regression settings.
Discussion and Limitations.
The performance gain on rel-stack is limited, as discussed in Section 5.1. We also observe divergent results across tasks in the rel-trial dataset. A potential cause might be the inherent limitation of the prediction head for the entity regression task rather than the GNN model itself, as noted in the original RelBench paper (Robinson et al., 2024). In particular, study-adverse involves estimating an unbounded target (the number of patients with severe outcomes), while site-success predicts a bounded success rate. One possible explanation is that unbounded predictions may be more challenging given the inherently constrained design of the regression head. Enhancing prediction heads for regression tasks is an important direction for future work.
@@ 5.3 Recommendation
Experimental Setup.
Recommendation tasks involve predicting a ranked list of top $K$ target entities for a given source entity at a specific seed time, where $K$ is predefined per task. This requires computing pairwise scores between source and target entities. We follow RelBench implementation and use two types of prediction heads: a two-tower GNN  (Wang et al., 2019a) and an identity-aware GNN (ID-GNN)  (You et al., 2021). The two-tower GNN calculates pairwise scores through the inner product of source and target node embeddings and is trained using the Bayesian Personalized Ranking loss  (Rendle et al., 2012). ID-GNN computes scores by applying an MLP to the embedding of the target entity within the subgraph sampled around each source entity and is trained using binary cross-entropy loss. Consistent with the original implementation, we use two-tower GNN for tasks in rel-amazon dataset and ID-GNN for tasks in the remaining datasets. The evaluation metric is Mean Average Precision (MAP) @$K$, where higher values indicate better performance. All results are averaged over five seeds.
Results.
Results are presented in Table 5.3. RelGNN  achieves better or same performance compared to baselines on all 9 tasks. These consistent improvements across diverse datasets and tasks underscore the versatility of our approach and demonstrate its effectiveness in modeling complex relational information and extracting key predictive signals crucial for recommendation.
Discussion and Limitations.
One limitation comes from the ID-GNN prediction head, which restricts recommendation candidates to nodes sampled within the source entity’s subgraph for efficiency. This limitation is quantified by the locality score (Yuan et al., 2024), which measures the fraction of ground-truth targets present in the sampled subgraph. A high locality score suggests users tend to re-engage with previously interacted entities, while a low score indicates a preference for novel items. ID-GNN struggles on low-locality tasks where ground-truth targets often fall outside the sampled subgraph. Accordingly, RelGNN shows greater gains on high-locality tasks (e.g., rel-stack) than on low-locality ones (e.g., rel-hm, rel-trial). Designing improved prediction heads for recommendation tasks is orthogonal to our core contribution and remains a valuable direction for future work.
@@ 6 Related Work
Several works have explored the use of GNNs for learning on relational data  (Schlichtkrull et al., 2018; Cvitkovic, 2019; Šír, 2021; Zahradník et al., 2023; Kanatsoulis et al., 2025). These works investigated different GNN architectures that utilize the relational structure. More recently, Fey et al. (2024) introduced Relational Deep Learning (RDL) (cf. Sec. 2.2), establishing a new subfield of machine learning. RDL has enabled various research opportunities, such as advancements in relational graph construction algorithms, GNN architectures, and task-specific prediction heads. Yuan et al. (2024) focused on improving recommendation tasks by addressing limitations of the currently employed two-tower and pair-wise prediction heads. In contrast, our work focuses on improving GNN architectures applied to all task types, offering an orthogonal contribution. Recently, a series of works have explored the use of large language models (LLMs) for heterogeneous graphs. HiGPT (Tang et al., 2024) introduced a language-enhanced heterogeneous graph tokenizer combined with LLMs. Wydmuch et al. (2024) proposed leveraging LLMs for predictive tasks in RDL, showing improvements on certain node-level tasks.
Meta-paths in Heterogeneous Graphs.
Meta-paths (Sun et al., 2011) are widely used in heterogeneous graph learning to capture semantic relationships between different types of entities (Shang et al., 2016; Dong et al., 2017; Hu et al., 2018; Shi et al., 2018; Wang et al., 2019b; Fu et al., 2020). A meta-path is a manually designed sequence of node and edge-types in a heterogeneous graph intended to capture specific relational patterns. For example, in an academic graph, a meta-path Author–Paper–Author models co-authorship, and a meta-path Author–Paper–Conference–Paper–Author captures co-conference patterns. While effective in many settings, meta-paths have some known limitations (Shi et al., 2016; Hu et al., 2020; Shi, 2022). They typically rely on manual specification, which requires domain expertise and may introduce bias or overlook important patterns. Designing an effective set of meta-paths for complex graphs can be time-consuming and may fail to comprehensively capture all relevant interactions.
@@
The literature of knowledge graphs  (Bordes et al., 2013; Wang et al., 2014, 2017) differs from RDL in terms of the tasks being tackled. Knowledge graph models mainly focus on completion tasks like predicting missing entities (e.g., Q: Who is the author of Harry Potter? A: J.K. Rowling) or missing relationships (e.g., Q: Did Yoshua Bengio win a Turing Award? A: Yes). In contrast, RDL focuses on making predictions about entities or groups of entities in the future timestamp (e.g., Will a customer churn in the next month? How much will a customer spend in the upcoming week?)